OpenStudy (anonymous):

Please helpp ! How many solutions does the system of equations have? 3. 2x=-10y+6 and x+5y=3 -one -two -three -infinittely many -none

OpenStudy (anonymous):

infinitely many

OpenStudy (anonymous):

Help With These Too Please ? (: How many solutions does the system of equations have? 4. y=5x+7 and 3y-15x=18 -one -two -inifinetly many -none 5. x-2y=6 and 3x-6y=18 -one -two -infinitely many 6. y-7x=-14 and 7y-49x=-2 -one -two -infinetily many -none